40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 6 00:14:10 2014, Jenne, Update, LSC, ALS is iffy tonight 
    Reply  Thu Aug 7 12:34:56 2014, Jenne, Update, LSC, Suspensions not kicking? 
Message ID: 10335     Entry time: Wed Aug 6 00:14:10 2014     Reply to this: 10345
Author: Jenne 
Type: Update 
Category: LSC 
Subject: ALS is iffy tonight 

The ALS system is iffy tonight.

After putting the cable back to the RF spectrum analyzer (it had been taken to test the frequency counter setup, and not put back), I had a good Yarm beatnote, but again this evening the Xarm beatnote is small.  I touched up the PSL table alignment (very, very little needed, but it did double my peak height).  I *think* that this is happening because we haven't settled into a good IFO alignment place, so the arm pointing keeps changing very slightly, which means that the PSL ALS alignment needs touching.  Anyhow, even after alignment the Xarm beatnote is only -36 dBm at 81 MHz.  It should be at least -25 dBm or so, although I haven't seen it any larger than about -35 dBm since the IFO beam was lost last Friday.

I am not able to hold ALS lock long enough to scan the arms and find the IR resonances.  The only optics that I am actuating on this evening are the 2 ETMs.  When I lose lock and look at the watchdogs, the ETMs are the only optics that have largeish numbers, which comes from the ALS lockloss.  So, I don't think I am suffering from the ITM suspension kicks tonight.  Rather, I think that it's that the ALS system isn't tuned up nicely.

I think that it is past time we tuned up and checked out the ALS PDH setup.  Q:  Can you please measure the loop TFs for both of the ALS PDH boxes tomorrow?  At the very least we want to know what we're working with. 

Evan:  What is the status with the ISS? 

I am going to try tomorrow to look at the suspensions, and see if I can track anything down.  I feel like I see the kicks more often when the arms are locked, i.e. we are sending an LSC signal to them.  The LSC POS signal is a factor of a few hundred larger than the damping SUSPOS signal is.  Are we saturating something somewhere?  Why is this a new thing?  We certainly do see kicks when the LSC is not engaged, so this may not be the right path, but it is something concrete to look at.

ELOG V3.1.3-